Assuming the lasting time of laptop batteries are normally distributed with a mean of 11 hours and a standard deviation of ) 0.7 hours.
a) If we randomly select 8 batteries, what is the probability that the average lasting time of these laptops is longer than 11.5 hours.
b) In a group of 30 laptops, approximately how many of them will last less than 10 hours?
c) Knowing that a lasting time of a laptop is 20% below the third quartile, how many hours can this laptop last?
1
Expert's answer
2020-12-24T17:41:04-0500

Let "X=" the lasting time of laptop batteries: "X\sim N(\mu, \sigma^2)"

Given "\mu=11h, \sigma=0.7h"

a) "\bar{X}\sim N(\mu,\sigma^2/n)"

Given "n=8"


"P(\bar{X}>11.5)=1-P(\bar{X}\leq11.5)"

"=1-P(Z\leq\dfrac{11.5-11}{0.7/\sqrt{8}})\approx1-P(Z\leq2.020305)"

"\approx1-0.978324=0.021676"

"P(\bar{X}>11.5)=0.021676"


b)


"P(X<10)=P(Z<\dfrac{10-11}{0.7})"

"\approx P(Z<-1.428571)\approx0.076564"

"30(0.076564)=2"


(c) The third quartile Q3 is the 75th percentile of a data set. 


"P(Z<\dfrac{X-11}{0.7})=0.75"

"\dfrac{X-11}{0.7}\approx0.674490"

"X\approx11+0.7(0.674490)=11.472143"

"\dfrac{80\%}{100\%}\cdot11.472143=9.18(h)"

9.18 hours